Given the following data set, answer the questions below: x 11 12 13 14 15 p(x) .20 .20 .30 .05 .25 1. What is the table above called? 2. What information is provided from the table? 3. What is the total probability of the sample space? 4. P(X > 13) 5. P(X ≥ 13) 6. P(X ≤ 14) 7. P(X = 15) 8. F(15) 9. F(14) 10. E(X) 11. V(X)
Answer #1

4.P(14)+P(15)=0.05+0.25=0.30

5. P(13)+P(14)+P(15)=0.3+.05+.25=0.6

6.P(15)=0.25

7.Same as sixth one

10. 0.2*11+0.2*12+0.3*13+0.05*14+0.25*15

=12.95

3.total probability of sample space=0.2+0.2+0.3+0.05+0.25=1